OpenStudy (anonymous):

A particle travels in the xy-plane so that the position is given by x(t) = cos(t) and y(t) = e-2t - 1. What is the total distance traveled by the particle for 0 ≤ t ≤ 4?

OpenStudy (anonymous):

is y(t)= e- 2 t-1

OpenStudy (anonymous):

or e^(-2t) -1

OpenStudy (anonymous):

y(t) = e^(-2t) ...sorry about that

OpenStudy (anonymous):

\[\int_{0}^{4}\sqrt{x^2+y^2}dt\]

OpenStudy (anonymous):

\[\sqrt{cos^2(t)+(e^{-2t}-1)^2}\]

OpenStudy (anonymous):

so how do i integrate this now?

OpenStudy (anonymous):

good question, I wouldn't know how to do it by hand but this being physics I guess you can use computer

OpenStudy (anonymous):

oh right...ok so let me do it on teh calculator n see what i get

OpenStudy (anonymous):

i thought since it was a calculus question, it would involeve some type og integrating of velocity, or accelaration, something along those lines

OpenStudy (anonymous):

i got 4.664 as my final answer

OpenStudy (anonymous):

try posting it in math section too because I gtg

OpenStudy (anonymous):

ok thanks for the help =)